# Matrix decompositions in Quantum Optics: Takagi/Autonne, Bloch-Messiah/Euler, Iwasawa, and Williamson

### 07 Mar 2024 v1

@@ I Introduction
In Sec. III we discuss some basic facts about the symplectic group which plays an important role later on. The discussion presented borrows from many excellent references such as Serafini Serafini (2017), Arvind Arvind et al. (1995) and Nicacio Nicacio (2021).
@@
In Sec. VI we discuss the pre-Iwasawa and Iwasawa decompositions. In Ref. Iwasawa (1949), K. Iwasawa introduced the decomposition that bears his name for general semisimple Lie groups. In this manuscript we obtain this decomposition for the specific case of the symplectic group. Both of these find applications in the study of group-theoretic properties of the symplectic group (see for example Arvind Arvind et al. (1995), Prop. 2.29 of Gosson De Gosson (2006), Page 2 of Haberman and Haberman Habermann and Habermann (2006) and Page 179 of Folland Folland (1989)). Moreover, the Iwasawa decomposition finds applications in e.g., obtaining keyrate bound in quantum key distribution Lodewyck and Grangier (2007) and discussing properties of CV cluster states González-Arciniegas et al. (2021).
@@ II.1 The spectral theorem and matrix square roots
Note that we use the term a matrix square root and not the matrix square root since one can pick the plus or minus sign for any of the eigenvalues appearing in the last equation. Regardless of the choice of sign, note that any matrix square root satisfies $\sqrt{\bm{A}}\sqrt{\bm{A}}=\bm{A}$. In the special case where all the eigenvalues are non-negative (positive), the unitarily diagonalizable matrix is termed positive semi-definite PSD (or positive definite PD) and then its unique PSD (PD) square root is obtained by always taking the plus signs in the equation above. The spectral theorem also applies to real-symmetric matrices where in Eq. (1) the unitaries are replaced by orthogonal matrices and $a_{i}\in\mathbb{R}$. Similarly, one can construct real PSD matrices as the symmetric matrices that have non-negative eigenvalues, these too have unique square roots, just like in the hermitian case. These conclusions, namely the replacement of unitary matrices by orthogonal matrices when discussing real-symmetric instead complex-hermitian matrices extend to the polar decomposition and SVD discussed below.
@@ II.2 The polar decomposition
II.3 Singular value decomposition
Once the polar decomposition of a matrix is at hand, one can further build a singular value decomposition by unitarily diagonalizing the PSD matrix $\bm{P}$ in Eq. (3) as
@@ II.3 Singular value decomposition
Since both $\bm{U}^{\dagger}$ and $\bm{W}$ are unitary, their product is unitary and thus we can identify a unitary matrix (or its transpose, conjugate or conjugate-transpose) as $\bm{V}^{\dagger}=\bm{U}^{\dagger}\bm{W}$ ($\bm{V}^{*}=\bm{U}^{\dagger}\bm{W}$, $\bm{V}^{T}=\bm{U}^{\dagger}\bm{W}$ or $\bm{V}=\bm{U}^{\dagger}\bm{W}$ respectively) to write the singular-value decomposition of $\bm{A}$ as
@@ IV Takagi/Autonne decomposition
where $\bm{W}$ is unitary, $\bm{W}\bm{W}^{\dagger}=\mathbb{1}_{\ell}$ and $\lambda_{i}\geq 0$. This decomposition is a singular value decomposition (SVD), albeit one with a special symmetry that makes it explicit that the matrix being decomposed is symmetric; indeed the decomposition on the right-hand side of the Eq. (52) makes it explicit that the object on the left-hand side is symmetric.
@@ V Bloch-Messiah/Euler Decomposition
Since $\bm{N}$ is PD, so is $\bm{N}^{2}$ and as such it has a unique square-root. Therefore, by knowing the Tagaki/Autonne decomposition of $\bm{M}$ we can obtain the eigendecomposition of $\bm{N}$
@@ VI Pre-Iwasawa and Iwasawa Decompositions
The Iwasawa decomposition gives the decomposition of a symplectic matrix in terms of three matrices from different subgroups of the symplectic group Arvind et al. (1995). Two of these groups already appeared before, namely the group of diagonal symplectic group (which is Abelian) and the compact subgroup which is isomorphic to the unitary group. The third subgroup that we will need for the Iwasawa decomposition is the nilpotent symplectic subgroup $\mathrm{N}(\ell)$. The matrices from this subgroup have the form
@@
The equality on the right hand side is derived by using the properties of the blocks (cf. Eq. (III)) as follows
@@ VII Williamson Decomposition
Williamson’s theorem states that given a real symmetric positive definite matrix $\bm{\Sigma}\in\mathbb{R}^{2\ell\times 2\ell}$ ($\bm{\Sigma}^{T}=\bm{\Sigma}>0$) there exists a real symplectic matrix $\bm{S}\in\mathrm{Sp}(2\ell,\mathbb{R})$ such that
@@
We follow the steps outlined in Idel et al. (2016, 2017) and define
@@
Numerically, directly obtaining the eigendecomposition of $\bm{\Psi}^{-2}$ does not always lead to the correct orthogonal matrix satisfying Eq. (176), as each eigenvalue of $\bm{\Psi}^{-2}$ is pair-wise degenerate. In particular they can always be swapped which can lead to an incorrect symplectic form in the right-hand side of Eq. (176), where not all the elements above the diagonal are non-negative.

### 13 Mar 2024 v2

@@ I Introduction
In Sec. III we discuss some basic facts about the symplectic group which plays an important role later on. The discussion presented borrows from many excellent references such as Serafini Serafini (2017), Arvind et al. Arvind et al. (1995) and Nicacio Nicacio (2021).
@@
In Sec. VI we discuss the pre-Iwasawa and Iwasawa decompositions. In Ref. Iwasawa (1949), K. Iwasawa introduced the decomposition that bears his name for general semisimple Lie groups. We provide details of this decomposition for the specific case of the symplectic group. Both of these find applications in the study of group-theoretic properties of the symplectic group (see for example Arvind Arvind et al. (1995), Prop. 2.29 of Gosson De Gosson (2006), Page 2 of Haberman and Haberman Habermann and Habermann (2006) and Page 179 of Folland Folland (1989)). Moreover, the Iwasawa decomposition finds applications in e.g., obtaining keyrate bounds in quantum key distribution Lodewyck and Grangier (2007) and discussing properties of CV cluster states González-Arciniegas et al. (2021).
@@ II.1 The spectral theorem and matrix square roots
Note that we use the term a matrix square root and not the matrix square root since one can pick the plus or minus sign for any of the eigenvalues appearing in the last equation. Regardless of the choice of sign, note that any matrix square root satisfies $\sqrt{\bm{A}}\sqrt{\bm{A}}=\bm{A}$. In the special case where all the eigenvalues are non-negative (positive), the unitarily diagonalizable matrix is termed positive semi-definite PSD (or positive definite PD) and then its unique PSD (PD) square root is obtained by always taking the plus signs in the equation above. The spectral theorem also applies to real-symmetric matrices where in Eq. (1) the unitaries are replaced by orthogonal matrices and $a_{i}\in\mathbb{R}$. Similarly, one can construct real PSD matrices as the symmetric matrices that have non-negative eigenvalues, these too have unique square roots, just like in the hermitian case. These conclusions, namely the replacement of unitary matrices by orthogonal matrices when discussing real-symmetric instead complex-hermitian matrices extend to the polar and singular-value decompositions discussed below.
@@ II.2 The polar decomposition
II.3 Singular-value decomposition
Once the polar decomposition of a matrix is at hand, one can further build a singular-value decomposition (SVD) by unitarily diagonalizing the PSD matrix $\bm{P}$ in Eq. (3) as
@@ II.3 Singular-value decomposition
Since both $\bm{U}^{\dagger}$ and $\bm{W}$ are unitary, their product is unitary and thus we can identify a unitary matrix (or its transpose, conjugate or conjugate-transpose) as $\bm{V}^{\dagger}=\bm{U}^{\dagger}\bm{W}$ ($\bm{V}^{*}=\bm{U}^{\dagger}\bm{W}$, $\bm{V}^{T}=\bm{U}^{\dagger}\bm{W}$ or $\bm{V}=\bm{U}^{\dagger}\bm{W}$ respectively) to write the SVD of $\bm{A}$ as
@@ IV Takagi/Autonne decomposition
where $\bm{W}$ is unitary, $\bm{W}\bm{W}^{\dagger}=\mathbb{1}_{\ell}$ and $\lambda_{i}\geq 0$. This decomposition is an SVD, albeit one with a special symmetry that makes it explicit that the matrix being decomposed is symmetric; indeed the decomposition on the right-hand side of the Eq. (52) makes it explicit that the object on the left-hand side is symmetric.
@@ V Bloch-Messiah/Euler Decomposition
Since $\bm{N}$ is PD, so is $\bm{N}^{2}$ and as such it has a unique square-root. Therefore, by knowing the Takagi/Autonne decomposition of $\bm{M}$ we can obtain the eigendecomposition of $\bm{N}$
@@ VI Pre-Iwasawa and Iwasawa Decompositions
The Iwasawa decomposition gives the decomposition of a symplectic matrix in terms of three matrices from different subgroups of the symplectic group Arvind et al. (1995). Two of these groups have already appeared before, namely the group of diagonal symplectic matrices (which is Abelian) and the compact subgroup $\mathrm{C}(\ell)$ which is isomorphic to the unitary group $\mathrm{U}(\ell)$. The third subgroup that we will need for the Iwasawa decomposition is the nilpotent symplectic subgroup $\mathrm{N}(\ell)$. The matrices from this subgroup have the form
@@
The equality on the right-hand side is derived by using the properties of the blocks (cf. Eq. (III)) as follows
@@ VII Williamson Decomposition
Williamson’s theorem states that given a real symmetric PD matrix $\bm{\Sigma}\in\mathbb{R}^{2\ell\times 2\ell}$ ($\bm{\Sigma}^{T}=\bm{\Sigma}>0$) there exists a real symplectic matrix $\bm{S}\in\mathrm{Sp}(2\ell,\mathbb{R})$ such that
@@
We follow Idel et al.Idel et al. (2016, 2017) and define
@@
Numerically, directly obtaining the eigendecomposition of $\bm{\Psi}^{-2}$ does not always lead to the correct orthogonal matrix satisfying Eq. (176), as each eigenvalue of $\bm{\Psi}^{-2}$ is pair-wise degenerate. In particular they can always be swapped which can lead to an incorrect symplectic form on the right-hand side of Eq. (176), where not all the elements above the diagonal are non-negative.